### v2

# 1]OpenDataLab 2]Shanghai Artificial Intelligence Laboratory

Abstract
Supervised Fine-Tuning (SFT) Large Language Models (LLMs) fundamentally rely on high-quality training data. While data selection and data synthesis are two common strategies to improve data quality, existing approaches often face limitations in static dataset curation that fail to adapt to evolving model capabilities. In this paper, we introduce Middo, a self-evolving Model-informed dynamic data optimization framework that uses model-aware data selection and context-preserving data refinement. Unlike conventional one-off filtering/synthesis methods, our framework establishes a closed-loop optimization system: (1) A self-referential diagnostic module proactively identifies suboptimal samples through tri-axial model signals-loss patterns (complexity), embedding cluster dynamics (diversity), and self-alignment scores (quality); (2) An adaptive optimization engine then transforms suboptimal samples into pedagogically valuable training points while preserving semantic integrity; (3) This optimization process continuously evolves with the model’s capability through dynamic learning principles. Experiments on multiple benchmarks demonstrate that our Middo consistently enhances the quality of seed data and boosts LLMs’ performance, improving accuracy by $7.15\%$ on average while maintaining the original dataset scale. This work establishes a new paradigm for sustainable LLM training through dynamic human-AI co-evolution of data and models.
[…]
1 Introduction
Large Language Models (LLMs) have revolutionized artificial intelligence by achieving state-of-the-art performance across diverse domains, from natural language understanding [67, 17] to mathematical reasoning [7, 16] and code generation [6, 2]. This success is largely attributed to Supervised Fine-Tuning (SFT), where models undergo rigorous training on high-quality, human-aligned datasets to ensure outputs closely match human expectations. Crucially, the quality of these datasets directly dictates the model’s ultimate capabilities: noisy or suboptimal training data can lead to degraded performance, while meticulously curated data unlocks advanced reasoning, generalization, and robustness. As LLMs scale, the adage “garbage in, garbage out” becomes increasingly important—highlighting the urgent need for systematic methods to optimize training data quality.
Existing approaches primarily fall into two categories to improve data quality: data selection [4, 65, 31, 19, 66, 32, 28] and data synthesis [10, 51, 40, 57, 35, 14]. Data selection methods filter raw datasets using heuristic rules (e.g., length filters) [61] or statistical metrics like perplexity (PPL) [35] and Instruction-Following Difficulty (IFD) [30] to retain “high-quality” samples. Conversely, data synthesis leverages advanced LLMs (e.g., GPT-4 [1]) to generate new training examples, often through prompting or distillation [33]. While both strategies improve data quality, they suffer from critical limitations. Selection methods are typically static, applying fixed criteria that ignore the evolving needs of the model during training. Similarly, synthesis approaches often discard original data, wasting potentially valuable information, and risk generating distributionally narrow or redundant examples. These one-time data curation methods fail to adaptively refine data along with the model’s progress.
Figure 1: Comparison of different dataset and different models before and after Middo optimization.
To overcome these limitations, we propose Middo, Model-informed Dynamic Data Optimization, a self-evolving framework that unifies model-aware data selection with context-preserving data refinement. Unlike static approaches, Middo establishes a closed-loop optimization system where data curation dynamically adapts to the model’s evolving capabilities. The framework operates through three core mechanisms: (1) A self-referential diagnostic module that proactively identifies suboptimal training samples using three model signals: loss patterns (to detect complexity mismatches between data and model proficiency), embedding cluster dynamics (to assess diversity gaps in the latent space), and self-alignment scores (to evaluate data quality against the model’s own knowledge). (2) An adaptive optimization engine that transforms these suboptimal samples into pedagogically valuable training points. For example, overly complex samples may be simplified through stepwise decomposition, while low-diversity clusters are enriched with controlled extension—all while preserving the original data’s semantic intent. (3) A dynamic principle that iteratively updates the training dataset based on the model’s progress, ensuring that data difficulty and diversity scale with the model’s capabilities. By integrating these components, Middo not only maximizes the utility of existing data but also bridges the gap between static data curation and adaptive model training.
Experiments across multiple benchmarks demonstrate Middo ’s effectiveness especially on low-quality datasets. Models trained with Middo optimized data achieve consistent performance gains over baselines, improving accuracy by $7.15\%$ on average while maintaining the original dataset scale. Notably, Middo-trained models exhibit stronger abilities to address hard problems, solving more than three times the number of challenging test problems (e.g., MATH, GPQA) compared to models trained on static datasets. These results validate that sustainable LLM advancement requires co-evolving data and models—a paradigm shift from today’s disjointed curation practices.
2 Related Work
2.1 Synthetic Data Generation
Synthetic data generation is a key technique for augmenting LLM fine-tuning. Early methods [13, 54] introduce perturbation-based approaches to enhance data diversity, using character-level [3] and word-level [53] modifications. These methods rely on fixed transformation rules, limiting adaptability.
LLMs have been leveraged for scalable data synthesis [48, 21, 10, 51, 40, 57, 35, 27]. Self-instruct methods [52] generate instruction-response pairs, while Evol-Instruct [55] and Auto-Evol-Instruct [60] refine data complexity iteratively. However, these methods remain static, failing to adapt as models improve. Recent approaches integrate model feedback into data generation [22, 39, 34, 33], incorporating student model signals for adaptive synthesis. LLM2LLM [26] is an iterative data augmentation strategy that enhances low-data fine-tuning by using a teacher LLM to generate synthetic training data from incorrect student LLM predictions and I-SHEEP [41] uses an iterative self-enhancement paradigm.
2.2 Data Selection
Data selection is crucial for LLM fine-tuning, as high-quality and informative data directly impacts model performance [64, 56]. Early heuristic-based methods rely on surface-level statistics like item frequency [44] and repetition count [25], but they also lack adaptability to model evolution.
Recent work explores LLM-driven data selection, optimizing for quality, diversity, and complexity [4, 65, 31, 19, 66, 32, 28, 11, 23]. The IFD metric [30] enables models to self-select training instances by comparing loss with and without the instruction, while other methods [58, 8, 38, 62] use LLM self-assessment for efficiency. Further advancements integrate LLM-based evaluation mechanisms. AlpaGasus [5] and LIFT [56] use structured prompts for data assessment, while DEITA [36] introduces a multi-dimensional scoring system based on complexity and quality.
3 Methodology
An overview of our Middo is shown in Figure 2. We first introduce the overall pipeline of Middo in Section 3.1, then elaborate on the three core components: complexity optimization (Section 3.2), diversity optimization (Section 3.3), and quality optimization (Section 3.4).
Figure 2: The Middo pipeline: a closed-loop, iterative dynamic optimization framework for LLM fine-tuning. It comprises three core modules that leverage model feedback: Loss Patterns identify overly complex samples, which are then simplified; Self-alignment Scores evaluate data quality, transforming low-quality samples into high-quality ones; and Embedding Cluster Dynamics detect sparse data points and expand the data distribution through targeted augmentation. Middo ensure the training set continually evolves to better align with the model’s capabilities.
3.1 Middo Pipeline
As depicted in Figure 2, our Middo framework establishes an iterative data-model co-evolution loop driven by tri-axial signal analysis, along with three interconnected data optimization mechanisms, each targeting distinct dimensions of training sample selection: (1) Loss patterns, to identify samples with mismatched complexity (overly challenging) relative to the current model’s capability through loss trajectory analysis. (2) Embedding cluster dynamics, to detect coverage gaps in the semantic space, ensuring balanced conceptual representation. (3) Self-alignment scores, for quality filtering to leverage the model’s self-evaluation capacity to flag low-confidence or inconsistent responses through automated alignment scoring.
At each iteration, these parallel signal analyzers jointly select suboptimal samples, which are then regenerated through context-aware synthesis—preserving original semantic intent while enhancing pedagogical value. The refined dataset immediately feeds back into model training, creating a dynamic feedback loop where improved model capabilities inform subsequent optimization cycles. Notably, the optimized dataset remains similar in data size, without extending large data synthesis, leading to an efficient data optimization. This self-referential mechanism ensures continuous alignment between data characteristics and model evolution. The following sections systematically elaborate on the implementation of each signal-specific optimization module and their synergistic integration.
3.2 Loss Patterns: Complexity Optimization
Complexity Selection.
Complexity reflects the “difficulty” or “compositionality” of data. A good dataset usually requires a smooth complexity distribution of data for training [15, 47]. Therefore, we introduce Loss Patterns, which targets overly challenging samples by modifying them to maintain a balanced and learnable training set [62]. During fine-tuning, the loss for a sample $(X_{i},Y_{i})$ is computed as the likelihood of predicting successive tokens given the instruction $X_{i}$ and its context. We denote the loss before and after training as $\mathcal{L}_{\text{pre}}(X_{i},Y_{i})$ and $\mathcal{L}_{\text{post}}(X_{i},Y_{i})$, respectively.
Intuitively, we consider both the loss before and after training to select the complex data. Specifically, we classify samples based on their loss evolution: samples with both low $\mathcal{L}_{\text{pre}}$ and $\mathcal{L}_{\text{post}}$ are considered easy, while those with high values in both metrics remain difficult, indicating excessive complexity. A sample is included in the complex subset $\mathcal{D}^{\text{hard}}$ if its $\mathcal{L}_{\text{pre}}$ and $\mathcal{L}_{\text{post}}$ both exceed the thresholds $\tau_{\text{pre}}$ and $\tau_{\text{post}}$, respectively. For adaptive refinement, the thresholds are dynamically computed. See Appendix 8.3 for details on the dynamic threshold settings used throughout the paper.
Complexity Optimization.
For complex data optimization, instead of discarding difficult samples, we transform them into simpler, more manageable forms. Specifically, we replace samples in $\mathcal{D}^{\text{hard}}$ with their simplified counterparts, $\mathcal{D}^{\text{hard}^{\prime}}$. This is achieved by an automatic process in which a LLM analyzes and summarizes the complex instructions [60], then simplifies them step by step while preserving the core educational content. An example is shown in Appendix Figure 11. This iterative transformation process updates the dataset by replacing overly complex samples with refined versions that offer more effective training samples. As training continues, this adaptive approach ensures a continuous alignment between data complexity and model capability.
3.3 Embedding Cluster Dynamics: Diversity Optimization
Diversity Selection.
Diversity is crucial for ensuring broad concept coverage and a uniform data distribution. Embedding Cluster Dynamics identifies sparse data points that signal underrepresented regions in the dataset. We extract sentence embeddings from the last hidden layer $\mathcal{H}^{(L)}$ of the model trained in the previous iteration, using average pooling, then compute the cosine similarity between each data point and find the $k$-nearest neighbors $\mathcal{N}_{k}(X_{i})$ for each data $X_{i}$. A lower average cosine similarity among these neighbors $\mathcal{N}_{k}(X_{i})$ indicates the data is positioned in a sparser region. Thus, the data points whose average cosine similarity score (diversity score) is below a threshold are selected for optimization.
Diversity Optimization.
To enhance diversity-balanced distribution, we augment the sparse subset $\mathcal{D}^{\text{sparse}}$ by incorporating examples from their corresponding $\mathcal{N}_{k}(X_{i})$ as demonstrations to generate new samples. This process generates an expanded set $\mathcal{D}^{\text{sparse}{\prime}}$, which is then integrated back into the dataset. An instance can be found in Appendix Figure 13. This structured augmentation strategy ensures that the data distribution becomes both broader and more balanced, ultimately improving the model’s generalization.
3.4 Self-alignment Scores: Quality Optimization
Quality Selection.
High-quality data is essential for fine-tuning, as poor-quality samples can degrade performance [64]. To reduce manual annotation costs, many approaches use the LLM-as-a-Judge paradigm [5, 56]. To achieve this, instead of relying on an external judge, we leverage the fine-tuned model itself to assess data quality via Self-alignment Scores, effectively incorporating the model’s own feedback. Specifically, for each instruction-response pair $(X_{i},Y_{i})$ in $\mathcal{D}$, the model generate scores $\mathcal{S}{\pi}^{\text{ins.}}(X_{i})$ for instruction and $\mathcal{S}{\pi}^{\text{res.}}(X_{i},Y_{i})$ for instruction-response pair based on three key metrics $\pi$ from AlignBench [37]: Clarity, Completeness, and Factuality. The final quality score $\mathcal{S}(X_{i},Y_{i})$ is obtained by averaging these scores. These samples with scores below a similar dynamic threshold are identified as low-quality, forming the seed dataset $\mathcal{D}^{\text{low}}$.
Quality Optimization.
To refine $\mathcal{D}^{\text{low}}$, we use LLMs to automatically analyze and improve these samples via tailored evolution strategies (prompt templates and examples are provided in the Appendix Figure 12). This process converts low-quality samples into higher-quality versions, denoted as $\mathcal{D}^{\text{low}\prime}$. The dataset is then updated by replacing the original low-quality samples with $\mathcal{D}^{\text{low}\prime}$, maintaining the dataset size while progressively enhancing its overall quality.
In each iteration, after the three data selection and optimization processes described above, the optimized dataset is then fed back for the next round of model training.
4 Experiment
4.1 Settings
Data Optimization Configurations.
We conduct optimization on the Alpaca [49] and WizardLM [55] datasets. For a fair comparison, we also include a rewritten version of Alpaca, where responses are generated by GPT-4o-mini, in our optimization process. Each dataset undergoes three iterations of optimization. Demonstrating that our method does not require a powerful external model, we synthesize data using DataDreamer [42] with GPT-4o-mini, setting both temperature and top_p to 1.0 to ensure diversity. A detailed analysis of the computational cost is provided in Appendix 7, and the effects of the number of neighbors and iteration counts are discussed in Appendix 8.
Training and Evaluation Settings.
We fine-tune LLaMA-3.1-8B [12] and Mistral-7B-v0.3 [20] using LLaMA-Factory [63] with the specific hyperparameters detailed in Appendix 9.5. For each iteration of Middo’s optimization, the base model is fine-tuned for one epoch on the dataset optimized in that specific iteration to mitigate the risk of overfitting to the data [26, 34]. Evaluation is conducted using OpenCompass [9], with vLLM [24] for acceleration. To validate the effectiveness and generalization capabilities of our approach, we assess model capabilities in general knowledge using IFEval [67] and MMLU [17]; mathematical problem-solving on GSM8K [7] and MATH [16]; code generation on HumanEval [6] and MBPP [2]; and commonsense reasoning on Hellaswag [59] and GPQA [45].
4.2 Main Results
The evaluation results on all benchmarks over various data iterations and models are presented in Table 1. We can see that Middo consistently enhances model performance across all benchmarks, achieving an average accuracy increase of $7.15\%$ over three iterations on the Alpaca dataset based on LLaMA-3.1-8B, all while preserving the original data scale. Moreover, when extending our experiments to Mistral-7B-v0.3, we observed an average improvement of $4.75\%$, further underscoring the robustness and adaptability of our framework across different model architectures.
Setting General Math Code Reasoning Average
MMLU IFEval GSM8K MATH HumanEval MBPP Hellaswag GPQA
Base Model: LLaMA-3.1-8B
Alpaca init 47.46 41.09 35.63 4.96 39.63 37.40 48.11 5.56 32.48
iter1 50.13 45.77 43.67 10.62 40.24 39.20 56.37 13.64 37.45
iter2 41.82 44.63 50.11 12.40 39.63 41.40 59.22 18.18 38.42
iter3 51.32 43.20 51.18 12.92 39.63 41.80 58.78 16.67 39.63
Alpaca 4o-mini init 32.82 44.04 57.09 17.78 51.22 45.20 53.70 24.24 40.76
iter1 41.09 43.47 54.21 17.34 51.22 46.00 59.11 21.72 41.77
iter2 44.69 47.96 57.62 18.50 52.44 45.40 57.37 19.70 42.96
iter3 38.58 48.11 58.68 18.30 46.95 46.80 52.37 28.79 42.32
Wizard init 46.12 46.14 53.30 12.72 40.24 48.00 53.05 12.12 38.96
iter1 48.39 50.11 54.44 13.80 46.95 45.00 63.54 20.20 42.80
iter2 48.86 49.48 55.12 13.90 48.78 45.20 58.63 18.18 42.29
iter3 47.18 50.79 54.51 11.70 43.29 45.40 62.97 20.20 42.01
Base Model: Mistral-7B-v0.3
Alpaca init 27.66 43.22 22.21 3.88 29.27 28.80 44.17 0.51 24.97
iter1 31.31 45.62 29.57 5.82 30.49 33.80 42.73 14.65 29.25
iter2 26.87 49.46 31.69 6.84 31.71 31.00 53.95 5.56 29.64
iter3 38.73 44.01 34.80 6.64 26.22 31.40 44.86 11.11 29.72
Alpaca 4o-mini init 31.56 43.14 44.88 9.64 42.07 37.80 46.25 21.21 34.56
iter1 31.33 47.93 45.19 8.72 37.20 41.32 41.32 19.70 34.09
iter2 28.83 47.92 48.90 11.34 35.37 38.40 42.63 27.27 35.08
iter3 28.96 50.78 48.60 10.10 32.32 39.00 32.95 20.20 32.86
Wizard init 40.71 50.95 44.96 8.10 35.98 35.60 53.98 9.09 34.92
iter1 41.39 51.18 44.43 9.44 37.80 38.60 59.01 17.17 37.38
iter2 33.87 51.71 47.08 9.26 39.02 38.40 66.18 19.70 38.15
iter3 43.26 49.80 41.09 10.02 41.46 34.60 66.02 22.22 38.56
Table 1: Performance comparison on different benchmarks using LLaMA-3.1-8B and Mistral-7B-v0.3 as base models. We use Alpaca, Alpaca-4o-mini, and Wizard as the optimization datasets for Middo. The init means training on the original dataset, while iter means training on the optimized dataset. Both init and iter settings are trained for one epoch. The best performance on average is highlighted in bold and the second best is underlined.
On the Alpaca dataset, the average score increased progressively with each iteration. Across the MMLU, GSM8K, MATH, and MBPP benchmarks, we observed consistent, step-by-step improvements over multiple iterations. This showcases the versatility of our approach, which excels in general capabilities, mathematics, and coding. Notably, accuracy on GSM8K improved by $15.55\%$, and Hellaswag saw an $11.11\%$ increase when evaluated on the LLaMA-3.1-8B model. For Mistral-7B-v0.3, we observed an $11.07\%$ improvement on MMLU, a $12.59\%$ increase on GSM8K, and a $10.6\%$ gain on GPQA. These results underscore the effectiveness of our method in driving performance gains and highlight the cumulative benefit of our iterative optimization process.
Further Validation on 4o-mini Rewritten Data.
Steady improvements observed on the 4o-mini rewritten Alpaca dataset—averaging a $2.2\%$ increase overall, with MMLU showing an impressive $11.87\%$ boost—demonstrate that these gains are not merely a result of using 4o-mini data. This illustrates that our framework intrinsically enhances data quality and model performance. Importantly, we achieve these improvements without resorting to stronger variants such as GPT-4o [18], reinforcing the robustness and general applicability of our method.
Initial Dataset Quality.
Our experiments reveal that higher-quality datasets require fewer modifications to reach optimal performance. On LLaMA-3.1-8B, for instance, while the Alpaca dataset achieves peak performance at third iteration, the 4o-mini rewritten Alpaca required only two iterations, and the Wizard dataset reaches its best performance in just one round.
Comparison with Other Works.
Method Size General Math Code Reasoning Average
MMLU IFEval GSM8K MATH HumanEval MBPP Hellaswag GPQA
Alpaca 52.0kk 47.46 41.09 35.63 4.96 39.63 37.40 48.11 5.56 32.48
Data Selection
Alpaca-clean 51.7kk 47.21 43.92 43.90 4.20 29.27 43.40 60.17 5.56 34.70
Superfiltering 7.8kk 39.96 37.80 44.50 5.38 40.85 44.00 42.38 27.27 35.27
Superfiltering GPT4 7.8kk 37.71 34.35 53.68 11.00 9.15 45.60 57.81 2.53 31.48
Long 1.0kk 25.51 14.75 56.33 16.56 13.41 45.60 25.83 0.00* 24.75
AlpaGasus 9.2kk 33.98 48.82 43.82 6.06 35.98 42.40 44.50 18.18 34.22
Data Augmentation
I-SHEEP 8.4kk 23.61 29.61 43.14 8.28 32.32 32.60 41.83 0.00* 26.42
Alpaca-GPT4 52.0kk 51.94 38.68 50.87 10.28 17.07 43.60 63.02 0.51 34.50
WizardLM 70.0kk 46.12 46.14 53.30 12.72 40.24 48.00 53.05 12.12 38.96
Middo Alpaca 57.6kk 51.32 43.20 51.18 12.92 39.63 41.80 58.78 16.67 39.63
MiddOnly Alpaca 8.8kk 43.47 40.78 65.20 15.58 51.83 47.60 58.65 17.68 42.60
Middo Alpaca-4o-mini 63.1kk 44.69 47.96 57.62 18.50 52.44 45.40 57.37 19.70 42.96
MiddOnly Alpaca-4o-mini 24.9kk 41.50 45.66 60.80 20.06 46.34 48.00 55.01 24.75 42.77
Table 2: Results of Middo compared to other baseline methods. The best and second best results are highlighted in bold and underlined, respectively. Our method outperforms all baselines in the average score. *Note that 0.00 indicates that the method did not solve any examples. $\dagger$ Denotes training solely on Middo-generated data.
We compare Middo with both data selection (Alpaca-clean[46], Superfiltering [29], Long [61], AlpaGasus [5]) and data augmentation (Alpaca-GPT4 [43], I-SHEEP [34], WizardLM [55]) methods on the Alpaca dataset.
We use the optimal dataset obtained through Middo from Alpaca for comparison with other baselines. Additionally, to ensure a relatively fair comparison with data selection methods, we include a dataset that only uses the optimized data without incorporating any unoptimized samples, referred to as MiddOnly, to isolate the effect of the optimization process and make a direct comparison with data selection approaches.
Results in Table 2 show our method achieves the highest average score of $42.96$, outperforming all other approaches. Notably, even when using only the optimized subset MiddOnly Alpaca, our method delivers a robust average score of $42.60$. This demonstrates that iterative improvement is not primarily driven by data size, but rather by the effectiveness of our dynamic data selection and optimization process in identifying and generating data with high learning value for models.
5 Analysis
5.1 Ablation Studies
To assess the effectiveness of Middo and the contribution of each optimization pipeline, we conduct ablation experiments with the LLaMA-3.1-8B model on the Alpaca dataset. Specifically, we analyze the following ablations: (a) w/o loss: removes Loss Patterns. (b) w/o neighbor: excludes Embedding Cluster Dynamics. (c) w/o score: removes Self-alignment Scores.
Table 3: Ablation study on the development set. We report the performance of the model with different ablations. The ablations include removing the loss patterns, embedding cluster dynamics and self-alignment scores separately. The best performance is highlighted in bold.
Iter. Ablations IFEval MATH HumanEval Hellaswag Avg.
iter1 w 45.77 10.62 40.24 56.37 38.25
w/o loss 42.49 10.11 39.02 59.53 37.79
w/o neighbor 39.01 10.82 42.07 57.86 37.45
w/o score 43.48 10.20 36.59 48.40 34.67
iter2 w 44.63 12.40 39.63 59.22 38.97
w/o loss 42.28 9.92 42.68 58.21 38.27
w/o neighbor 46.75 10.26 34.76 46.66 34.61
w/o score 44.18 11.76 39.02 51.38 36.58
iter3 w 44.24 12.92 39.63 59.25 39.01
w/o loss 43.18 12.42 36.59 55.30 36.87
w/o neighbor 40.12 12.46 34.15 56.83 35.89
w/o score 45.17 7.92 40.85 54.67 37.15
The ablation results in Table 3 consistently show that removing any part of the framework leads to a decline in performance across multiple iterations, reinforcing that each component plays a significant role in the overall performance. This trend holds across the second (iter2) and third (iter3) iterations, where the removal of any pipeline consistently results in suboptimal performance, further highlighting the importance of balancing complexity, diversity, and quality in the optimization process. These findings underscore the necessity of the full framework for achieving optimal results.
5.2 Effect of Selected Data Scale
We investigate the impact of the different scales of the selected and optimized data in this section by varying the thresholds for data selection. Results are illustrated in Figure 3. We observe that increasing the size of the refined data initially leads to an upward trend in performance; however, once the refined data exceeds a certain threshold, performance begins to decline. To maintain the potential for further iterative improvement, we set the refined data size at a moderate level that optimally balances the cost and benefit of the optimization process. In the first iteration, each component selects approximately $5\%$ of the data for refinement. By controlling the parameter $m$, the amount of data refined can adaptively change as the model’s capability increases. Detailed data sizes selected in each iteration are provided in Appendix 11.
5.3 Data Analysis
Dynamic Iterative Improvement.
For a deeper understanding of how Middo transforms the dataset, we provide an analysis of its impact on data complexity, diversity, and quality.
Complexity.
To quantify how Middo modulates dataset complexity, we analyze the loss distribution evolution through optimization cycles. As shown in Figure 4, the original dataset exhibits a long-tailed distribution with extreme loss values up to $12.99$. After applying Middo, the maximum loss decreases by $71.05\%$ to $3.76$, indicating successful mitigation of overly complex samples and the distribution mode shifts leftward, suggesting better alignment between data complexity and model capability. This transformation demonstrates our framework’s ability to adaptively prune pathological samples while preserving pedagogically valuable challenges.
Diversity.
To analyze the diversity of the dataset after applying Middo, we visualize the data distribution using t-SNE [50]. Figure 5 reveals how the augmented data points are distributed relative to the original data. Notably, most of the augmented samples are located at the peripheries of the clusters, effectively filling in the sparsely populated regions. This distribution indicates that Middo is not merely adding redundant data but is instead enhancing the overall coverage of the latent space. By strategically augmenting the dataset at the cluster edges, Middo improves the diversity and ensures a more uniform distribution of data points, ultimately contributing to better model generalization.
Figure 3: Performance comparison of Middo on the Alpaca dataset with varying refined data sizes. The x-axis represents the number and percentage of data selected for refinement, while the y-axis shows the average accuracy across three iterations. To ensure fairness, we guarantee that the data after refinement is the same.
Quality.
The self-alignment score trajectories across different iterations are presented in Figure 6. The observed trend indicates a gradual increase in the average score as the iterations progress. This improvement signifies that the quality of the data is becoming more closely aligned with the model’s evolving capabilities. Through the adversarial self-play mechanisms and iterative quality refinement, the model is able to assess and enhance the quality of both the instructions and responses within the dataset. As the self-alignment scores increase, it reflects that the refined data is not only more accurate but also more consistent with the model’s internal standards and expectations. This detailed evolution of the self-alignment scores provides critical insights into the dynamic process of dataset optimization, confirming that our approach effectively transforms low-quality samples into high-quality learning material over successive iterations.
Figure 4: Loss distribution comparison before and after applying Middo. The density curve reflects the relative frequency of data points within specific loss intervals. The inset subfigure highlights the maximum loss reduction from $12.99$ to $3.76$.
Figure 5: t-SNE visualization of the Alpaca dataset before and after applying Middo. The original dataset is shown in light blue, while the augmented data is in dark blue. The dark blue points tend to occupy the sparsely populated regions of the light blue point distribution.
Figure 6: Self-alignment score evolution across iterations. The x-axis represents the number of iterations, while the y-axis shows the average self-alignment score.
6 Conclusion
In this paper, we present Middo, a model-informed dynamic data optimization framework that transforms LLM fine-tuning via closed-loop learning. Unlike traditional static methods, Middo establishes a self-evolving system that continuously adapts to the model’s evolving capabilities. It employs three core mechanisms: complexity optimization refines overly complex samples using loss patterns, ensuring the training data remains appropriately challenging; diversity optimization enhances dataset diversity by analyzing embedding cluster dynamics; and quality optimization leverages self-alignment scores to evaluate and improve the quality of training samples. Experiments on multiple benchmarks demonstrate that Middo consistently boosts LLMs’ performance, achieving an average accuracy improvement of $7.15\%$ while maintaining the original data scale on LLaMA-3.1-8B. Ablation studies confirm the effectiveness of each component, underscoring the importance of balancing complexity, diversity, and quality. Middo ’s adaptability and model-awareness make it a powerful tool for sustainable LLM training. Moreover, our approach paves the way for future research in adaptive training that continuously optimizes learning efficiency.
Limitations
Despite its promising results, Middo has several limitations: (1) Middo relies on the model being fine-tuned itself for identifying data quality and complexity. This means that the approach requires a sufficiently capable base model, and the performance may be limited if the base model is not strong enough to generate meaningful diagnostics for data refinement. (2) Middo does not currently utilize Reinforcement Learning, which could further enhance data refinement, especially for complex or subjective tasks. (3) The closed-loop optimization system may lead to higher computational costs as the dataset grows or updates become more frequent, presenting scalability challenges. (4) Middo may propagate biases present in the initial training data, limiting fairness and generalization if the base model is trained on biased data. These limitations highlight areas for future improvement, such as integrating RL, optimizing for scalability, and addressing data biases.
Acknowledgments
This work is supported by Shanghai Artificial Intelligence Laboratory. Zinan Tang is an intern at Shanghai Artificial Intelligence Laboratory.